Sally travels 200m west 490m south find the displacement and distance?
1
Expert's answer
2020-03-10T11:28:45-0400

The distance


"d=200+490=690\:\rm m."

The displacement


"s=\sqrt{200^2+490^2}=529\:\rm m."